# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v2

@@ 4.2. Harmonia: RL Formulation
State. Table 1 summarizes the state features for Harmonia’s two RL agents. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics when making their decisions. The state features are represented as a 7-dimensional tuple (observation vector) for both agents:
@@
The observation vector size directly impacts Harmonia’s storage and computational overheads (see §4.6), so we (1) carefully select a limited set of highly relevant features, and (2) group state feature values into a small number of bins (Table 1) based on workload characterization. For example, request size can fall into eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB) requiring three bits to represent. In total, we use 32 bits for encoding the state representation.
@@
$req\_type$ Request type (read/write) 2 1
$req\_size$ Request size (in pages) 8 3
$acc\_intr$
@@
Action. The action space of Harmonia’s agents depends on the number of devices in the HSS. The data-placement agent receives the state features of incoming write requests, and its action is to determine the best-fit device to place request data. The data-migration agent continuously looks at the state features of previously-placed pages to determine their target device for migration. Therefore, the two agents do not perform their actions in synchronization. In an HSS with two devices, (1) the data-placement agent can place data in the fast or the slow device, and (2) the data-migration agent can migrate data to the fast or the slow device. The action sets of both agents are small and extensible, resulting in less computational complexity and easier multi-device scalability. We encode actions with 4 bits, allowing extensibility to more storage devices.
@@
where $L_{t}$ is I/O request latency at time step $t$. The reward is higher (lower) if the data is placed in the fast (slow) device. We design a simple reward for the data-placement agent because its objective is to minimize the latency of the ongoing I/O request.
@@ 4.5. Design of Harmonia’s RL Agents
Experience Buffer. Experiences are tuples of $\langle State,Action,\\ Reward,NextState\rangle$ collected by the inference thread of each agent and stored in an experience buffer to use as training data. We use separate experience buffers for the two agents because they have different goals and experiences. Based on empirical analysis, we set the experience buffer size to store the most recent 1000 experiences. A larger experience buffer can capture more access patterns but comes with significant storage overhead. The experience buffers reside in the host DRAM. Harmonia uses experience replay (dqn, ; singh2022sibyl, ), where a batch of experiences is randomly sampled from the experience buffer and used as training data for the RL agent.
Training and Inference networks. We separate the RL agent’s training and inference networks to allow parallel execution of training and inference, and remove training latency overhead from I/O request handling. Training network weights are periodically copied to the inference network (after every 1000 requests) for the inference network to adapt to current workload and HSS conditions. For the training and inference networks of both agents, we use an identical simple feed-forward neural network structure (singh2022sibyl, ) with one hidden layer of 10 neurons (de1993backpropagation, ) that uses use the swish activation function (ramachandran2017searching, ). Harmonia’s data-placement and data-migration policies select the action with the maximum predicted Q-value.
@@
Convergence of RL Agents in Harmonia. Figs 7(a) and  7(b) show the convergence of Harmonia’s data-placement and data-migration policies in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 4) on a performance-optimized HSS.
@@
We make two key observations: (1) The placement policy converges in less than 6000 I/O requests without prior training. (2) The migration policy takes longer to converge (around 9000 I/O requests) due to indirect and delayed reward to the migration agent. The convergence of Harmonia’s policies indicates that the agents learn to coordinate their actions and make optimal decisions.
@@ 4.6. Overhead Analysis
In this section, we provide an analysis of Harmonia’s overhead in terms of (1) inference and training latencies, (2) storage overhead.
@@
Harmonia has low inference and training latencies because of its (1) lightweight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). Harmonia’s two RL agents have identical networks. We perform training and inference on host CPU as the neural networks are lightweight, and their weights fit in the CPU on-chip caches in our evaluated system.
@@
4.6.2. Storage Overhead.
Harmonia’s storage overhead includes (1) RL agents’ neural networks, (2) experience buffers, (3) migration queue, (4) migration agent’s reward, and (5) address mapping information. First, neural network weights are represented using a half-precision floating-point format. With 90 16-bit weights, each network requires $\sim$1.5 KiB of memory. The four networks (training and inference networks for two agents) take a negligible 6 KiB of memory, so we store their weights in the CPU caches. Second, to train each RL agent, we use an experience buffer located in the DRAM that consumes 100 KiB to store 1000 experiences. Hence, the total storage overhead of two experience buffers is 200 KiB. Overall, Harmonia’s RL agents require 206 KiB of memory. Third, the migration queue holds the logical block address (32 bits) and target device ID (4 bits) for each queue entry. The migration queue size is 10 in our experiments, requiring 50 bytes of DRAM. Fourth, we store the latencies of 50 incoming I/O requests temporarily to assign a delayed reward to the migration agent, which consumes  100B. Fifth, we maintain address mapping information for HSS devices (tsukada2021metadata, ). Harmonia requires 32 bits to store the state features for each page (see Table 1). This overhead is less than $\sim$0.8of the total storage capacity when using a 4-KiB data placement granularity.
@@ 6.1. Performance Analysis
Average I/O Request Latency and End-to-End Throughput. Figs. 8(a) and  8(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance- and cost-optimized HSS configurations. Figs. 9(a) and 9(b) show their end-to-end throughput on performance- and cost-optimized HSS, respectively.
@@
Tail Latency. Figs. 10(a) and 10(b) show the 99th and 99.99th percentile I/O latencies (tail latency) of Harmonia and baselines on performance-optimized HSS.
@@
Performance Sensitivity to Migration Queue Size. Fig.  11(a) shows the impact of migration queue size on HSS performance.
@@
Performance Sensitivity to Migration Agent’s Reward. Fig.  11(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3).
@@ 6.2. Multi-Programmed Workloads
Figs. 13(a) and 13(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 5) on performance- and cost-optimized HSS, respectively.
@@ 6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 3). We do not evaluate other baselines as they are designed for fixed HSS configurations and lack easy extensibility. Figs. 14(a) and 14(b) show the performance of Sibyl and Harmonia on a Tri- and Quad-HSS. We discuss Harmonia’s extensibility in §7.
@@
6.4. Lifetime of Devices in HSS
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to data originating from the workload. A higher WA adversely impacts device lifetimes. Figs. 15(a) and 15(b) show WA of Harmonia and baselines on performance- and cost-optimized HSS, respectively.
@@ 6.4. Lifetime of Devices in HSS
Write Traffic Distribution in Harmonia. Fig.  16 shows the write distribution of Harmonia’s migration policy across devices in performance- and cost-optimized HSS.
@@ 7. Discussion
Alternate Objectives and Reward Structures. Since Harmonia’s focus is to improve HSS performance, we design the RL agents’ reward structure based on I/O latencies. We can design Harmonia for other objectives such as device lifetimes, fairness, quality of service (QoS) and energy efficiency. For example, to improve device lifetimes, the state features and reward structure can include (1) writes performed on each device, (2) device endurance in terms of program/erase cycles, (3) write amplification (WA), and (4) device reliability properties (e.g., NAND flash type (SLC/TLC/QLC), ECC capability). To manage write traffic to each device, the reward should be appropriate to the device where the data is placed or migrated. Harmonia can accommodate fairness and QoS objectives by including factors such as bandwidth utilization, device queue depths and application requirements. We leave the exploration of these objectives and reward structures to future work.
Extensibility of Harmonia. Harmonia is easily extensible to a large number of storage devices with minimal designer effort. When a new device is added to the HSS, we only add (1) capacity utilization of the new device as a state feature (see Table 1), and (2) an output node in each RL agent’s neural network to represent the action of placing or migrating a page to the new device. However, scaling Harmonia to more devices can lead to two challenges: (1) the large state and action spaces can delay the convergence to an optimal policy, and (2) the metadata overhead increases.
Applicability to Other Systems. Harmonia can be adapted to other heterogeneous systems such as tiered storage, hybrid memory, disaggregated storage and networked storage, which have diverse constraints and requirements. In a tiered storage system, placement is always performed on the top-tier device, and hence Harmonia’s placement agent can be replaced by a heuristic approach. The migration agent gains significance due to data migration across tiers. We can add multiple migration agents for each pair of adjacent tiers. Hybrid memory systems have strict latency constraints, and require Harmonia to be redesigned for low-latency decision-making. Disaggregated storage systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) for large-scale memory pooling can leverage Harmonia’s benefits when deployed on CXL fabric controller. We hope Harmonia can assist in designing adaptive policies for other heterogeneous systems.
@@ 8. Related Work
In this section, we review other related work in HSS data management, and RL in storage/memory systems.
@@
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. Hot Random Off-loading (HRO) (lin2011hot, ) uses a 0-1 knapsack model to allocate or migrate files between the storage devices. Vengerov (vengerov2008reinforcement, ) uses a fuzzy rule base architecture to perform data migration in a hierarchical storage system. These works do not perform (1) real system evaluation or (2) combined optimization of placement and migration policies using multiple RL agents.
@@
RL in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) use RL to improve different aspects of storage systems. Prior works  (liu2019learning, ; wang2020reinforcement, ) use RL to improve cloud storage utilization. RLAlloc (li2023rlalloc, ) improves SSD throughput and Quality of Service (QoS). Other works (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ) use RL to optimize SSD maintenance activities (e.g., garbage collection) and error mitigation. Harmonia is orthogonal to all these works and can be used in conjunction with these techniques on an HSS.
@@
Multi-Agent RL Systems. Several works propose using multi-agent RL (MARL) for various system optimizations (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )). Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) use MARL to improve tail latency of Serverless Function-as-a-Service (FaaS), targeting different systems and objectives than Harmonia. Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) use MARL for dynamic partitioning of last level cache and multi-level caches. Shen et al. (shen2022multiagent, ) propose MARL-assisted cache cleaning for shingled managed recording drives. These techniques have four key differences from Harmonia: (1) their objectives are different from HSS data placement and migration, (2) they use identical RL agents unlike Harmonia’s heterogeneous agents, (3) they use joint actions and shared rewards for all agents, and (4) they use table-based RL unlike Harmonia’s neural network based RL. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.

### v3

@@ 4.2. Harmonia: RL Formulation
State. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics when making their decisions. Table 2 summarizes the state features for Harmonia’s two RL agents. Request type (req_typereq\_type) indicates a read or a write operation. Request size (req_sizereq\_size) is the number of 4KiB pages in the current I/O request. Access interval (acc_intracc\_intr) is the number of page accesses between two accesses to the same page, while access frequency (acc_freqacc\_freq) is the total number of accesses to the page. Fast-device capacity (fast_capfast\_cap) is the remaining free space in the fast storage device. Current device (curr_devcurr\_dev) indicates the storage device where the page currently resides. Migration interval (migr_intrmigr\_intr) is the number of migrations performed between two migrations of the current page. The state features are represented as a 7-dimensional tuple (observation vector) for both agents:
@@
The observation vector size directly impacts Harmonia’s storage and computational overheads (see §4.6), so we (1) carefully select a limited set of highly relevant features, and (2) group state feature values into a small number of bins (Table 2) based on workload characterization. For example, request size can fall into eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB) requiring three bits to represent. In total, we use 32 bits for encoding the state representation. Based on empirical analysis, we use additional bits for some state features (e.g., acc_intr, acc_freq, migr_intr) to represent the values accurately and provide extensibility to new workloads and more storage devices in the future.
Harmonia maintains a metadata table in the host DRAM to track the state features, and updates this table for each I/O request. Harmonia performs a metadata table lookup to observe the current state of a page in the HSS for placement and migration decisions.
@@
req_typereq\_type Request type (read/write) 2 1
req_sizereq\_size Request size (in pages) 8 3
acc_intracc\_intr
@@
Action. Harmonia’s agents operate independently and their action space depends on the number of devices in the HSS. The data-placement agent selects a device for each incoming write request, while the data-migration agent decides migration targets for existing data. In an HSS with two devices, the agents must choose between fast and slow devices. The small, extensible action sets reduce complexity and support multi-device scalability, with actions encoded with 4 bits to allow extension to more storage devices.
@@
where $L_{t}$ is I/O request latency at time step $t$. The reward is higher (lower) if the data is placed in the fast (slow) device. We design a simple reward for the data-placement agent because its objective is to minimize the latency of the ongoing I/O request. While placing data on the fast device yields lower latency and a higher reward, Harmonia does not always place data in the faster device for two reasons: (1) Fast device capacity is limited, and overusing the fast device can degrade performance due to factors such as queueing delays and garbage collection. (2) RL agents learn to maximize long-term cumulative reward. Hence, they learn to balance current latency against future access patterns and device utilization. Therefore, Harmonia places data on the slow device to prevent contention, and reserves the fast device for critical data.
@@ 4.3. Coordination Between Harmonia’s Agents
Impact of Coordinated vs. Uncoordinated Agents. To demonstrate the impact of coordination between Harmonia’s placement and migration agents, we compare the performance of Sibyl and Harmonia against Harmonia without coordinated agents in Figs. 6(a) and  6(b) on a performance-optimized and cost-optimized HSS, respectively. We design Harmonia without coordination by devising a reward for the migration agent that improves local performance (i.e., latency cost of page migrations) without considering the placement agent’s actions. While there are other possible reward structures, we choose one that does not facilitate coordination.
Figure 6. Performance of Sibyl, Harmonia without coordination and Harmonia on performance- (cost-) optimized HSS. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Harmonia without coordination by 35.1% (42.1%) on average on a performance- (cost-) optimized HSS. The two agents in Harmonia without coordination are oblivious to each other and take conflicting decisions. Second, Harmonia without coordination performs poorly in read-intensive workloads (e.g., RocksDB) because the migration agent does not prioritize migrating pages to the fast device that will likely be read in the future, which results in lower performance than Sibyl.
@@ 4.5. Design of Harmonia’s RL Agents
Experience Buffer. The two agents collect $\langle State,Action,Reward,NextState\rangle$ tuples (experiences) via their inference threads and store them in separate experience buffers, reflecting their distinct goals. The experience buffers reside in the host DRAM, and we use empirical analysis to set their sizes to store the most recent 1000 experiences. A larger experience buffer can capture more access patterns but increases storage overhead. Harmonia uses experience replay (dqn, ; singh2022sibyl, ) to sample a batch of experiences from the experience buffers to train its agents.
Training and Inference networks. We separate the RL agent’s training and inference networks to remove training latency from I/O request handling. Weights are copied from training network to inference network every 1000 requests to keep inference up to date. Both agents use identical simple feed-forward networks (singh2022sibyl, ) with one hidden layer of 10 neurons (de1993backpropagation, ) and swish activation (ramachandran2017searching, ). Actions are selected based on the highest predicted Q-value.
@@
Convergence of RL Agents in Harmonia. Figs 9(a) and  9(b) show the convergence of Harmonia’s data-placement and data-migration policies in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 5) on a performance-optimized HSS. Both agents start with no prior training (i.e., cold start).
@@
We make two key observations: (1) The placement policy converges in less than 6000 I/O requests. (2) The migration policy takes longer to converge (around 9000 I/O requests) due to indirect and delayed reward to the migration agent. The training loss remains low after convergence during the workload lifetime. After the initial convergence, Harmonia’s agents are well-equipped to adapt to new access patterns very quickly. They no longer need 6000 requests to start making good decisions when access patterns change.
Adaptability to Workload Changes. To demonstrate Harmonia’s adaptability to changes in workload access patterns, we design a synthetic benchmark using two workloads, LUN0 (write-intensive) and ResNet50 (read-intensive) (see Table 5), alternately every 5000 requests. Figs 10(a) and  10(b) show the access pattern variations in our benchmark in terms of virtual addresses accessed and request sizes. Fig 10(c) shows Harmonia’s convergence in terms of training loss for the first 20000 requests on a performance-optimized HSS. While the access patterns change significantly at every 5000 requests, each workload phase has considerable access pattern variations.
Figure 10. Workload access patterns represented in terms of (a) virtual addresses accessed, and (b) request sizes. Convergence of Harmonia shown in terms of training loss (c) for the first 20000 I/O requests. Lower is better.
Both agents start with no prior training (i.e., cold start). We make two key observations: (1) The rapid decline in training loss at the start indicates that both agents learn the optimal policy from a cold start in less than 5000 requests (Phase 1: LUN0). (2) When the access patterns change significantly after 5000 requests (Phase 2: ResNet50), the training loss remains low (nearly 0). (3) We observe a minor spike (shown in the inset plot) in training loss at every phase change. However, the training loss is still low, and Harmonia’s policies quickly converge again (in a few hundred I/O requests).
We conclude that Harmonia adapts quickly to abrupt changes in workload after the initial convergence.
@@ 4.6. Overhead Analysis
In this section, we analyse Harmonia’s overhead in terms of (1) storage overhead, and (2) inference and training latencies.
4.6.1. Storage Overhead.
[…]
@@
Harmonia has low inference and training latencies because of its (1) lightweight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). We perform training and inference on the host CPU as the neural networks are lightweight, and their weights fit in the CPU on-chip caches in our evaluated system.
@@
[…]
We conclude that Harmonia’s overheads are justified for the performance improvement it achieves for two key reasons. First, Harmonia’s RL agents have a significantly lower memory footprint compared to the that of best-performing prior approach, Sibyl. While Sibyl’s RL agent consists of 58 neurons, Harmonia’s RL agent is light-weight with only 19 neurons. Harmonia’s fewer neurons results in lower inference and training latency compared to Sibyl. Second, Harmonia provides a significant performance improvement (49.5% (31.7%) on performance- (cost-)optimized HSS) over Sibyl.
4.7. Architectural Differences over Prior Works
We describe Harmonia’s novelty and four key architectural differences compared to prior RL techniques (e.g., Sibyl) for HSS.
Separation of Objectives via Multi-Agent RL. Unlike Sibyl’s single-agent RL design, Harmonia decouples placement and migration into two specialized agents, which allows them to learn distinct policies: the placement agent focuses on immediate performance, while the migration agent optimizes long-term benefits. This design improves adaptability in dynamic workload and HSS conditions.
Proactive Data Migration. Harmonia learns to proactively migrate data (even in read-intensive workloads), rather than reacting only when the fast device capacity reaches 100% (as Sibyl does via eviction) or when there are updates from the application. The migration agent optimizes long-term HSS performance resulting in better, more robust handling of dynamic workloads.
Non-trivial RL Agent Formulation. Although Harmonia’s RL agent formulation shares some aspects (e.g., input features, latency-based reward) with Sibyl, it differs in three ways: (1) Harmonia’s neural networks are significantly more light-weight (90% fewer neurons), (2) The input features include migration characteristics (e.g., migration interval) to avoid ping-pong migrations, and (3) Harmonia’s reward structures are designed to optimize performance while facilitating coordination between the agents (See §4.3).
Extensibility. Harmonia’s multi-agent architecture is more amenable to further extensions such as integrating prefetching or load balancing agents.
@@ 6.1. Performance Analysis
Average I/O Request Latency and End-to-End Throughput. Figs. 11(a) and  11(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance- and cost-optimized HSS configurations. Figs. 12(a) and 12(b) show their end-to-end throughput on performance- and cost-optimized HSS, respectively. Performance and throughput are shown in terms of average request latency and IOPS, respectively, normalized to that of Fast-Only.
@@
Tail Latency. Figs. 13(a) and 13(b) show the 99th and 99.99th percentile I/O latencies (tail latency) of Harmonia and baselines on performance-optimized HSS.
@@
Performance Sensitivity to Migration Queue Size. Fig.  14(a) shows the impact of migration queue size on HSS performance.
@@
Performance Sensitivity to Migration Agent’s Reward. Fig.  14(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3).
@@ 6.2. Multi-Programmed Workloads
Figs. 16(a) and 16(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 6) on performance- and cost-optimized HSS, respectively.
@@ 6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 4). We do not evaluate other baselines as they are designed for fixed HSS configurations and lack easy extensibility. Figs. 17(a) and 17(b) show the performance of Sibyl and Harmonia on a Tri- and Quad-HSS. We discuss Harmonia’s extensibility in §7.
@@
6.4. Impact on Device Lifetimes
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to data originating from the workload. A higher WA adversely impacts device lifetimes. Figs. 18(a) and 18(b) show WA of Harmonia and baselines on performance- and cost-optimized HSS, respectively.
@@ 6.4. Impact on Device Lifetimes
Write Traffic Distribution in Harmonia. Fig.  19 shows the write distribution of Harmonia’s migration policy across devices in performance- and cost-optimized HSS.
@@
6.5. Performance on Unseen Workloads
We evaluate Harmonia’s performance on five unseen data-intensive workloads that are not used for hyperparameter tuning. We choose these workloads (wong2024baleen, ; berg2020cachelib, ; pan2021facebook, ) that use a flash cache in front of a bulk storage system from Meta’s production environment. We compare Harmonia to the best-performing prior approach, Sibyl, in Figs.  20(a) and 20(b) on a performance- (cost-) optimized HSS.
Figure 20. Performance of Sibyl and Harmonia on performance-optimized (top) and cost-optimized (bottom) HSS. Performance is shown in terms of average request latency normalized to Fast-Only. Lower is better.
We make two key observations. Harmonia outperforms Sibyl by 31.2% (31.5%) on average on performance- (cost-) optimized HSS. Second, while Harmonia’s hyperparameters may not be optimized for these workloads, its online-learning-based data-placement and data-migration policies adapt to changing workload access patterns better than Sibyl. We conclude that Harmonia’s generality and multi-agent RL framework provide consistent performance benefits across all unseen workloads.
@@ 7. Discussion
Alternate Objectives and Reward Structures. Harmonia focuses on improving HSS performance using RL agents with I/O latency-based rewards. However, it can be adapted for other objectives such as device lifetime, fairness, quality of service (QoS), or energy efficiency. For example, optimizing device lifetime may involve using write counts, endurance, write amplification, and device reliability (e.g., NAND type, ECC) in the input features and reward structures. Fairness and QoS can be addressed by incorporating metrics such as bandwidth, queue depth, and application requirements. We leave the exploration of these objectives and rewards to future work.
Extensibility of Harmonia. Harmonia is easily extensible to a large number of storage devices with minimal designer effort. Adding a new device to the HSS requires adding (1) its capacity utilization to the state features, which results in an additional input node per RL agent neural network, and (2) one output node per RL agent neural network. First, the capacity utilization of the new device consumes 3 bits of storage overhead (see Table 2). Second, the additional neural network nodes (1) add 1.3 KiB of storage overhead, and (2) increase inference and training latency by 30–40 $ns$. While Harmonia outperforms prior approaches, scaling to more devices leads to two challenges: (1) larger state and action spaces can delay policy convergence, and (2) an increase in metadata overhead.
Applicability to Other Systems. Harmonia can be extended to other heterogeneous systems such as tiered storage, hybrid memory, disaggregated storage and networked storage. In tiered storage, placement can be heuristic (always place in the top tier), while migration becomes key, potentially requiring multiple agents for each pair of adjacent tiers. Hybrid memory systems may require Harmonia to be redesigned for low-latency decision-making. Disaggregated systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) can benefit from deploying Harmonia on the CXL fabric controller. We hope Harmonia can aid in designing adaptive policies across diverse architectures.
@@ 8. Related Work
To our knowledge, Harmonia is the first work to use multi-agent RL for combined optimization of data-placement and data-migration policies in HSS. In this section, we review other related work in HSS data management, and RL in storage/memory systems.
@@
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. The Linux kernel offers utilities with heuristic solutions to manage multiple storage drives. For example, mdadm provides software RAID in the kernel, combining physical disks into a logical volume, but the burden of using mdadm for data placement and migration lies with applications. Similarly, LVM enables virtual partitions across multiple drives, allowing system administrators to resize them and move data across them. In contrast, Harmonia offers adaptive, application-transparent data placement and migration. For data migration, Vengerov (vengerov2008reinforcement, ) applies fuzzy rules in hierarchical storage and Hot Random Off-loading (lin2011hot, ) uses a 0-1 knapsack model. However, neither (1) conducts real-system evaluation or (2) jointly optimizes placement and migration with multiple RL agents.
@@
RL in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) apply RL to storage systems, focusing on cloud utilization (liu2019learning, ; wang2020reinforcement, ), SSD throughput and QoS (li2023rlalloc, ), and maintenance tasks such as garbage collection and error mitigation (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ). Harmonia is orthogonal to these works and can be integrated with them in HSS.
RL for Architectural Optimizations. Prior works use RL for data prefetching (e.g., (pythia, ; peled2015semantic, )), memory scheduling (e.g., (ipek2008self, ; multi_scheduler_HPCA_2012, )), cache replacement strategies (liu2020imitation, ) and network-on-chip arbitration (rl_NOC_AIDArc_2018, ; lin2020deepNOC, ). These works provide evidence that RL holds significant potential for intelligent high-performance computer systems.
@@
Multi-Agent RL Systems. Several works (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) use multi-agent RL (MARL) for system optimization. For example, Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) target FaaS latency, Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) optimize CPU caches, and Shen et al. (shen2022multiagent, ) address cache cleaning in SMR drives. These works use identical RL agents, joint actions, shared rewards, and table-based RL unlike Harmonia’s heterogeneous, autonomous, neural network-based agents. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.